Determine the root of f(x)=sinx+cos(1+x^2)-1 where x is in radians. Use 4 iterations of the secant method with initial guesses of xi-1=1.5 and xi=2.5


1
Expert's answer
2021-04-12T01:53:01-0400

Initialization. Two initial guesses "x_0" and "x_1" are chosen.

Iteration. For "n=1,2,3,..."


"x_{n+1}=x_n-f(x_n)\cdot\dfrac{x_n-x_{n-1}}{f(x_n)-f(x_{n-1})}"

Given "f(x)=\sin(x)+\cos(1+x^2)-1"

"x_0=1.5, x_1=2.5"



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 Step & x_0 & x_1 & x_2 & f(x_2) \\ \hline
 1 & 1.500000 & 2.500000 & 2.356929 & 0.669 842 \\
 \hdashline
 2 & 2.500000 & 2.356929 & 2.547287 & -0.082828 \\
 \hdashline
 3 & 2.356929 & 2.547287 & 2.526339 & 0.031471 \\
 \hdashline
 4 & 2.547287 & 2.526339 & 2.532107 & 0.000570 \\
 \hdashline
 5 & 2.526339 & 2.532107 & 2.532213 & -0.000004
\end{array}"

Root is : "x=2.532213"